# Retrograde predominance of small saturnian moons reiterates a recent retrograde collisional disruption

### v1

@@ Abstract
We find that in the inclination range 157–172 degrees the ratio of smaller moons (diameters less than 4 km) to larger moons (diameters greater than 4 km) is significantly larger than that of any other inclination range in the retrogrades. We denote this subset of the Norse group as the Mundilfari subgroup after its largest member. The incredibly steep slope of the Mundilfari subgroup’s size distribution, with a differential power law index of q = 6, points to this subgroup being created by a recent catastrophic collision proposed in Ashton et al. (2021).
@@ 1 Introduction
All four giant planets in the Solar System have moons that are thought to be planetesimals that were captured by their host planet during the final stages of planet formation. These moons, known as irregular moons, can be distinguished from their regular counterparts (moons that formed around their host planet as the planet was forming) by their orbits. The semi-major axes of irregular moons are typically much larger than those of regular moons, hundreds of planetary radii versus a few to tens of planetary radii. Additionally, irregular moons have moderate to large orbital eccentricities and a variety of inclinations, whereas the regular moons have near-circular orbits that are situated around the equator of the host planet. See Jewitt & Sheppard (2005), Nicholson et al. (2008), and Denk et al. (2018) for review articles on irregular moons.
@@
The other successful survey at finding saturnian irregulars used the Subaru Telescope between 2004 and 2007. Overall, this search was able to discover and track 45 new moons. Although there is currently no publication that describes this survey, both the reviews of Nicholson et al. (2008) and Denk et al. (2018) contain discussions of 25 of the 45 moons. Combining Phoebe with the moons found in the two aforementioned surveys, a total of 58 saturnian irregulars were known by 2019.
The orbits of irregular moons can be significantly tilted relative to the orbital plane of the host planet, which causes large oscillations in both inclination and eccentricity due to solar perturbations. Therefore, the time-averaged orbital elements are a much better measure of an irregular moon’s orbit than the osculating orbital elements. The time-averaged orbital elements of irregular moons can be calculated using N-body simulations (see Turrini et al. (2008); Jacobson et al. (2022)). Unless stated otherwise, anytime semi-major axis ($a$), eccentricity ($e$), or inclination ($i$) of an irregular moon is mentioned hereafter we are referring to the time-averaged value.
@@
The $10^{\circ}$ separation between the Gallic and Inuit groups is a factor of two smaller than the up to 20 variation in inclination that a saturnian irregular can experience (Turrini et al., 2008; Jacobson et al., 2022), which shows why calculating time-averaged orbital elements is vital for studying the orbital distribution of irregular moons. A good example is Saturn LX, with a current inclination of 44.4 it was initially thought that the moon was a member of the Inuit group. However, when using the time-averaged inclination 38.6, it was evident that Saturn LX belongs to the Gallic group (Jacobson et al., 2022).
@@
This paper details the results of a search for irregular moons of Saturn using the Canada France Hawaii Telescope (CFHT) from 2019 to 2021, and an analysis of the saturnian irregular moon population. A summary of the observations and the search method can be found in section 2. The 64 new moons discovered as a result of the survey are described in section 3. Analysis of the orbital distribution is found in section 4. A look at the size distribution of groups/subgroups is presented in section 5. The case for the Mundilfari subgroup being a recent collisional family is detailed in section 6. Finally, we give concluding remarks in section 7.
@@ 2 Observations
In the first 4 nights of July, 2019 CFHT acquired images of one of two fields near Saturn (one east and one west of the planet, with offsets shown in Fig. 1). Each night contained 44 sequential wide-band (gri) images that were acquired over an approximately 3 hour period. The 44 205-sec images from a night were combined using the shift and stack technique to detect moving irregular moons of Saturn down to diameters, $D$, of approximately 3 km ($m_{w}=26.3$). A total of 120 objects moving at Saturn-like rates were detected, of which 45 were eventually linked to previously designated moons. This search and a debiased size distribution over the entire moon system is detailed in Ashton et al. (2021). At this time, even the orbital sense (i.e. retrograde versus direct) of most of the orbits were not known.
@@ 3 Detections
Over all images in our survey we were able to observe all of the 58 previously known saturnian irregulars except for one, Bestla. Bestla’s large orbital inclination and node location results in it spending a greater fraction of its time nearly due north or south of Saturn, which our fields did not cover (Fig. 1); because other moons of comparable inclinations were both recovered and newly detected, this must partly be a timing issue of where Bestla was at the time. All of the previously known moons now have 14-year observational arcs at the minimum [^1] Even though we did not observe Bestla, another team measured it in 2015, providing a 14-year observational arc., and thus possess very high-precision orbits. Our survey produced 64 new irregular moons of Saturn, details of which are provided below.
@@ 3.1 Finding new moons
In total, 64 new irregular moons of Saturn have been discovered using our data set. One of these, S/2019 S 1 whose very small semimajor axis results in being visible outside Saturn’s glare for only a small fraction of the time, was announced in 2021 (Ashton et al., 2022); the rest were announced in May 2023. Additionally, we were able to detect 50+ more moon candidates that we were unable to link together due to an insufficient number of detections.
Figure 1: On-sky plot of the two fields used for this survey relative to Saturn (grey rectangles). All measurements of 5 of the 64 new moons detected are shown (circles) along with the best fit orbit (dashed lines). BJG: This sub-sample illustrates the challenges of multi-year linking. (Note that the apparent non-closure of the orbits is due to the projected saturnocentric positions being observed from the moving Earth.)
@@ 3.2 Linking back to past surveys
Like this survey, the 2003-2007 survey using Subaru detected many moon candidates that were unable to be connected into a long enough arc to produce confirmed moons. The precision of our multi-year orbits of the new moons is good enough that the Minor Planet Center was able to link back the majority of our linked moons to these previous candidates. Of the 64 new moons, 42 were matched to candidates found by Subaru from 2004–2007, ranging from a single night of observations to seven nights spread over multiple oppositions. If a solar system object is detected on at least two different nights in a single opposition, then the year of the first night becomes the discovery year. As such, 33 of the new moons ultimately have 2004–2007 as their discovery year and have one of those years in their temporary designation (see footnote 6).
@@ 4.1 Size
to estimate each moon’s size, where $m_{\sun}$ is the Sun’s apparent magnitude, and $H$, $p$, and $r$ are the Solar System absolute magnitude, geometric albedo, and radius in km of the moon, respectively. For this equation to be valid, the values for $H$, $p$, and $m_{\sun}$ all need to be in the same filter. We will use the V-band filter for our size calculations since that is the filter the Minor Planet Center uses for their values for $H$ of irregular moons [^2] https://www.minorplanetcenter.net/iau/NatSats/NaturalSatellites.html. For $m_{V,\sun}$ we use a value of -26.77, which comes from $M_{V,\sun}=4.80$ (Willmer, 2018)). Since only a few of the saturnian irregulars have measured albedos we have decided to go with a single value of $p_{V}=0.06$, which is the median value of the three saturnian irregular moons with measured albedos, $0.081\pm 0.02$ for Phoebe (Simonelli et al., 1999), $0.06\pm 0.03$ for Albiorix, and $0.050\pm 0.017$ for Siarnaq (Grav et al., 2015). Substituting in the values for $m_{V,\sun}$ and $p_{V}=0.06$, and turning $r$ into $D$ (diameter), Equation 1 becomes
@@
Table 1: A list of the irregular moons of Saturn with their semi-major axis ($a$), eccentricity ($e$), inclination ($i$), orbital period ($P$), V-band absolute magnitude ($H_{V}$), the group they belong to, and the subgroup we have assigned to them (if any). Orbital elements are taken from JPL [^5] https://ssd.jpl.nasa.gov/sats/elem/##\##refs and $H_{V}$ from MPC [^6] https://www.minorplanetcenter.net/iau/NatSats/NaturalSatellites.html. All inclinations are relative to the ecliptic except for Phoebe, which is relative to the Laplace plane.
Name/ $a$ $e$ $i$ $P$ H Group Subgroup
@@
Bebhionn (37) 0.1138 17.03 0.482 37.4 834.9 14.99 Gallic Albiorix
@@
Erriapus (28) 0.1170 17.51 0.462 38.7 871.1 13.71 Gallic Albiorix
Albiorix (26) 0.1092 16.33 0.47 38.9 783.5 11.17 Gallic Albiorix
@@
S/2019 S 6 0.1214 18.11 0.120 46.4 919.7 15.73 Inuit Siarnaq
S/2020 S 3 0.1207 18.05 0.144 46.1 908.0 16.38 Inuit Siarnaq
@@
Paaliaq (20) 0.1003 15.00 0.384 47.1 687.1 11.71 Inuit -
@@
S/2020 S 1 0.0758 11.34 0.337 48.2 451.1 15.92 Inuit Kiviuq
Siarnaq (29) 0.1195 17.88 0.311 48.2 895.9 10.61 Inuit Siarnaq
Kiviuq (24) 0.0756 11.31 0.182 48.9 449.1 12.67 Inuit Kiviuq
Ijiraq (22) 0.0758 11.34 0.353 49.2 451.5 13.27 Inuit Kiviuq
@@
Tarqeq (52) 0.1186 17.75 0.119 49.7 885.0 14.82 Inuit Siarnaq
Bestla (39) 0.1360 20.34 0.461 136.3 1087.5 14.61 Norse -
Narvi (31) 0.1289 19.29 0.449 143.7 1003.8 14.52 Norse -
@@
Skathi (27) 0.1041 15.58 0.265 149.7 728.1 14.41 Norse -
Hyrrokkin (44) 0.1226 18.34 0.331 150.3 931.9 14.34 Norse -
S/2019 S 19 0.1541 23.05 0.458 151.8 1318.0 16.51 Norse Kari
Kari (45) 0.1473 22.03 0.482 153 1231.0 14.49 Norse Kari
S/2004 S 21 0.1545 23.12 0.394 153.2 1325.4 16.21 Norse Kari
S/2004 S 36 0.1566 23.43 0.625 153.3 1352.9 16.11 Norse Kari
@@
Geirrod (66) 0.1488 22.26 0.539 154.4 1251.1 15.89 Norse Kari
@@
S/2006 S 1 0.1253 18.75 0.105 156.0 964.1 15.65 Norse Kari
S/2019 S 20 0.1583 23.68 0.354 156.1 1375.4 16.73 Norse Kari
@@
S/2019 S 15 0.1416 21.19 0.257 157.7 1161.5 16.59 Norse Mundilfari
Farbauti (40) 0.1356 20.29 0.248 157.7 1087.3 15.75 Norse Mundilfari
S/2004 S 37 0.1066 15.94 0.447 158.2 754.5 15.92 Norse Mundilfari
@@
Skoll (47) 0.1178 17.63 0.47 158.4 878.4 15.41 Norse Mundilfari
Bergelmir (38) 0.1288 19.27 0.144 158.7 1005.6 15.16 Norse Mundilfari
Thiazzi (63) 0.1576 23.58 0.511 158.8 1366.7 15.91 Norse Mundilfari
S/2019 S 5 0.1275 19.08 0.215 158.8 990.4 16.65 Norse Mundilfari
Beli (61) 0.1384 20.70 0.087 158.9 1121.8 16.09 Norse Mundilfari
S/2007 S 9 0.1348 20.17 0.36 159.3 1078.1 16.06 Norse Mundilfari
@@
Gunnlod (62) 0.1413 21.14 0.251 160.4 1158.0 15.57 Norse Mundilfari
@@
S/2020 S 7 0.1163 17.40 0.5 161.5 861.7 16.79 Norse Mundilfari
S/2020 S 9 0.1700 25.43 0.531 161.4 1535.0 16.02 Norse Mundilfari
@@
Table 2: Continuation of footnote 6.
Name/ $a$ $e$ $i$ $P$ H Group Subgroup
@@
S/2019 S 16 0.1555 23.27 0.25 162 1341.2 16.68 Norse Mundilfari
S/2006 S 13 0.1334 19.95 0.313 162 1060.6 16.05 Norse Mundilfari
S/2004 S 53 0.1556 23.28 0.24 162.6 1342.4 16.16 Norse Mundilfari
Jarnsaxa (50) 0.1289 19.28 0.219 163 1006.9 15.62 Norse Mundilfari
Gridr (54) 0.1287 19.25 0.187 163.9 1004.8 15.77 Norse Mundilfari
S/2019 S 10 0.1384 20.71 0.249 163.9 1123.0 16.66 Norse Mundilfari
S/2004 S 50 0.1494 22.35 0.45 164 1260.4 16.4 Norse Mundilfari
@@
Hati (43) 0.1317 19.70 0.375 164.1 1040.3 15.45 Norse Mundilfari
Fenrir (41) 0.1493 22.33 0.136 164.3 1260.3 15.89 Norse Mundilfari
@@
Eggther (59) 0.1326 19.84 0.157 165 1052.3 15.39 Norse Mundilfari
S/2004 S 52 0.1768 26.45 0.292 165.3 1634.0 16.5 Norse Mundilfari
S/2020 S 10 0.1692 25.31 0.295 165.6 1527.2 16.86 Norse Mundilfari
S/2004 S 41 0.1210 18.10 0.3 165.7 914.6 16.31 Norse Mundilfari
S/2004 S 42 0.1219 18.24 0.158 165.7 925.9 16.11 Norse Mundilfari
S/2004 S 39 0.1551 23.20 0.101 165.9 1336.2 16.14 Norse Mundilfari
@@
S/2006 S 14 0.1408 21.06 0.06 166.7 1152.7 16.5 Norse Mundilfari
Aegir (36) 0.1381 20.66 0.255 166.9 1119.3 15.51 Norse Mundilfari
@@
S/2020 S 6 0.1422 21.27 0.481 166.9 1168.9 16.55 Norse Mundilfari
S/2019 S 3 0.1142 17.08 0.249 166.9 837.7 16.22 Norse Mundilfari
S/2019 S 12 0.1397 20.90 0.476 167.1 1138.8 16.33 Norse Mundilfari
S/2004 S 44 0.1305 19.52 0.129 167.7 1026.2 15.82 Norse Mundilfari
S/2004 S 17 0.1317 19.70 0.162 167.9 1040.9 15.95 Norse Mundilfari
S/2004 S 28 0.1462 21.87 0.159 167.9 1220.7 15.77 Norse Mundilfari
Sat LXIV 0.1614 24.15 0.279 168.3 1420.8 16.15 Norse Mundilfari
Surtur (48) 0.1521 22.75 0.449 168.3 1296.5 15.77 Norse Mundilfari
Mundilfari (25) 0.1243 18.59 0.21 168.4 952.9 14.57 Norse Mundilfari
S/2006 S 17 0.1496 22.38 0.425 168.7 1264.6 16.01 Norse Mundilfari
S/2004 S 13 0.1233 18.45 0.265 169 942.6 16.25 Norse Mundilfari
@@
S/2006 S 18 0.1421 22.76 0.131 169.5 1298.4 16.1 Norse Mundilfari
Fornjot (42) 0.1667 24.94 0.214 169.5 1494.0 15.12 Norse Mundilfari
S/2019 S 4 0.1201 17.96 0.409 170.1 904.3 16.46 Norse Mundilfari
@@
S/2019 S 21 0.1767 26.44 0.155 171.9 1636.3 16.18 Norse Mundilfari
S/2019 S 8 0.1356 20.28 0.311 172.8 1088.7 16.28 Norse Phoebe
Sat LVIII 0.1745 26.10 0.148 172.9 1603.9 15.7 Norse Phoebe
S/2006 S 9 0.0963 14.41 0.248 173 647.9 16.48 Norse Phoebe
Ymir (19) 0.1535 22.96 0.337 173.1 1315.2 12.41 Norse Phoebe
@@
Greip (51) 0.1229 18.38 0.317 173.4 937.0 15.33 Norse Phoebe
S/2006 S 11 0.1318 19.71 0.144 174.1 1042.3 16.47 Norse Phoebe
S/2007 S 2 0.1066 15.94 0.232 174.1 754.9 15.59 Norse Phoebe
S/2019 S 7 0.1349 20.18 0.232 174.2 1080.3 16.29 Norse Phoebe
Gerd (57) 0.1400 20.95 0.517 174.4 1143.0 15.87 Norse Phoebe
Thrymr (30) 0.1359 20.33 0.467 174.8 1091.8 14.33 Norse Phoebe
Suttungr (23) 0.1296 19.39 0.116 175 1016.7 14.55 Norse Phoebe
@@
S/2019 S 13 0.1402 20.97 0.318 177.3 1144.9 16.68 Norse Phoebe
Angrboda (55) 0.1376 20.59 0.216 177.4 1114.1 16.17 Norse Phoebe
Alvaldi (65) 0.1471 22.00 0.238 177.4 1232.2 15.62 Norse Phoebe

### v2

@@ Abstract
We find that in the inclination range 157–172 degrees the ratio of smaller moons (diameters less than 4 km) to larger moons (diameters greater than 4 km) is significantly larger than that of any other inclination range in the retrogrades. We denote this subset of the Norse group as the Mundilfari subgroup after its largest member. The incredibly steep slope of the Mundilfari subgroup’s size distribution, with a differential power law index of q = 6, strengthens the hypothesis in Ashton et al. (2021) that this subgroup was created by a recent catastrophic collision, $10^{8}$ yr ago.
@@ 1 Introduction
All four giant planets in the Solar System have moons that are thought to be planetesimals that were captured by their host planet during the final stages of planet formation. These moons, known as irregular moons, can be distinguished from their regular counterparts (moons that formed around their host planet as the planet was forming) by their orbits. The semi-major axes of irregular moons are typically much larger than those of regular moons, hundreds of planetary radii versus a few to tens of planetary radii. Additionally, irregular moons have moderate to large orbital eccentricities and a variety of inclinations, whereas the regular moons have near-circular orbits that are situated around the equator of the host planet. See Jewitt & Sheppard (2005), Nicholson et al. (2008), and Denk et al. (2018) for review articles covering saturnian irregular moons.
@@
The other successful survey at finding saturnian irregulars used the Subaru Telescope between 2004 and 2007. Overall, this search was able to discover and track 45 new moons. This survey was eventually outlined in Sheppard et al. (2023). The reviews of Nicholson et al. (2008) and Denk et al. (2018) contain discussions of 25 of the 45 moons from this survey. Combining Phoebe with the moons found in the two aforementioned surveys, a total of 58 saturnian irregulars were known by 2019.
Strong solar perturbations on these distant moons cause periodic oscillations in both inclination and eccentricity, taking $\sim$500 yr for saturnians. Thus, time-averaged orbital elements are a more stable measure of an irregular moon’s orbit than the osculating orbital elements; these time-averaged orbital elements of irregular moons can be calculated using N-body simulations to monitor the evolution of each of the moon’s orbital elements over time (see Nesvorný et al. (2003); Turrini et al. (2008); Jacobson et al. (2022)). Simulations need to be integrated over many oscillations to obtain accurate time-averaged elements. Unless otherwise stated, anytime semi-major axis ($a$), eccentricity ($e$), or inclination ($i$) of an irregular moon is mentioned hereafter we refer to the time-averaged value.
@@
The $10^{\circ}$ separation between the Gallic and Inuit groups is a factor of two smaller than the up to 20 variation in inclination that a saturnian irregular can experience (Nesvorný et al., 2003; Turrini et al., 2008; Jacobson et al., 2022), which shows why calculating time-averaged orbital elements is vital for studying the orbital distribution of irregular moons. A good example is Saturn LX, with a current inclination of 44.4 it was initially thought that the moon was a member of the Inuit group. However, when using the time-averaged inclination 38.6, it was evident that Saturn LX belongs to the Gallic group (Jacobson et al., 2022).
@@
Although there is currently no consensus on the exact mechanism by which the initial population of irregular moons was captured, multiple theories have been proposed. The most popular are gas drag by an extended envelop and three-body interactions (either the planet capturing one member of a passing binary, or planetesimal capture during a planet-planet encounter). Refer to Nicholson et al. (2008) for a detailed discussion of the proposed capture scenarios. Knowing the orbital distribution of the initial irregular moon population would provide constraints on the capture mechanism. However, this is complicated by the fact that the irregular moon populations have likely undergone significant collisional alteration over the age of the Solar System (Bottke et al., 2010). Therefore, understanding the collisional history of an irregular moon system allows one to better constrain the initial population and thus the capture process, realising that only traces of the original captures are currently still present.
This paper details the results of a search for irregular moons of Saturn using the Canada-France-Hawaii Telescope (CFHT) from 2019 to 2021, and an analysis of the saturnian irregular moon population. A summary of the observations and the search method can be found in section 2. The 64 new moons discovered as a result of the survey are described in section 3. Analysis of the orbital distribution is found in section 4. A look at the size distribution of groups/subgroups is presented in section 5. The case for the Mundilfari subgroup being a recent collisional family is detailed in section 6. Finally, we give concluding remarks in section 7.
@@ 2 Observations
In the first 4 nights of July, 2019 CFHT acquired images of one of two fields near Saturn (one east and one west of the planet, with offsets shown in Fig. 1). Each night contained 44 sequential wide-band (gri) images that were acquired over an approximately 3 hour period. The 44 205-sec images from a night were combined using the shift and stack technique (Gladman et al., 1998; Kavelaars et al., 2004; Ashton et al., 2021) to detect moving irregular moons of Saturn down to diameters, $D$, of approximately 3 km ($m_{w}=26.3$). A total of 120 objects moving at Saturn-like rates were detected, of which 45 were eventually linked to previously designated moons. This search and a debiased size distribution over the entire moon system is detailed in Ashton et al. (2021). At this time, even the orbital sense (i.e. retrograde versus direct) of most of the orbits were not known.
@@ 3 Detections
Over all images in our survey we were able to observe all of the 58 previously known saturnian irregulars except for one, Bestla. Bestla’s large orbital inclination and node location resulted in it spending a large fraction of its time during these observations nearly due north or south of Saturn, which our fields did not cover (Fig. 1); because other moons of comparable inclinations were both recovered and newly detected, this must partly be a timing issue of where Bestla was at the time. All of the previously known moons now have 14-year observational arcs at the minimum [^1] Even though we did not observe Bestla, another team measured it in 2015, providing a 14-year observational arc., and thus possess very high-precision orbits. Our survey produced 64 new irregular moons of Saturn, details of which are provided below.
@@ 3.1 Finding new moons
In total, 64 new irregular moons of Saturn have been discovered using our data set. One of these, S/2019 S 1 whose very small semi-major axis results in being visible outside Saturn’s glare for only a small fraction of the time, was announced in 2021 (Ashton et al., 2022); the rest were announced in May 2023. Additionally, we were able to detect 50+ more moon candidates that we were unable to link together due to an insufficient number of detections.
Figure 1: On-sky plot of the two fields used for this survey relative to Saturn (grey rectangles). All measurements of 5 of the 64 new moons detected are shown (circles) along with the best fit orbit (dashed lines). This sub-sample illustrates the challenges of multi-year linking. (Note that the apparent non-closure of the orbits is due to the projected saturnocentric positions being observed from the moving Earth.)
@@ 3.2 Linking back to past surveys
Like this survey, the 2003-2007 survey using Subaru (Sheppard et al., 2023) detected many moon candidates that were unable to be connected into a long enough arc to produce confirmed moons. The precision of our multi-year orbits of the new moons is good enough that the Minor Planet Center was able to link back the majority of our linked moons to these previous candidates. Of the 64 new moons, 42 were matched to candidates found by Subaru from 2004–2007, ranging from a single night of observations to seven nights spread over multiple oppositions. If a solar system object is detected on at least two different nights in a single opposition, then the year of the first night becomes the discovery year. As such, 33 of the new moons ultimately have 2004–2007 as their discovery year and have one of those years as part of their temporary designation (see footnote 7).
@@ 4 Groups
For our analysis of the orbital distribution of the saturnian irregular moons we used the time-averaged orbital elements calculated by JPL [^2] https://ssd.jpl.nasa.gov/sats/elem/#\#refs. Their methods are detailed in Jacobson et al. (2022), where they simulated the orbits of the 58 irregular saturnian moons known prior to this work, plus S/2019 S 1, over a 5,000 year period. All inclinations are relative to the ecliptic except for Phoebe, which is relative to the Laplace plane.
@@ 4.1 Size
to estimate each moon’s size, where $m_{\sun}$ is the Sun’s apparent magnitude, and $H$, $p$, and $r$ are the Solar System absolute magnitude, geometric albedo, and radius in km of the moon, respectively. For this equation to be valid, the values for $H$, $p$, and $m_{\sun}$ all need to be in the same filter. We will use the V-band filter for our size calculations since that is the filter the Minor Planet Center uses for their values for $H$ of irregular moons [^3] https://www.minorplanetcenter.net/iau/NatSats/NaturalSatellites.html. For $m_{V,\sun}$ we use a value of -26.77, which comes from $M_{V,\sun}=4.80$ (Willmer, 2018). Since only a few of the saturnian irregulars have measured albedos we have decided to go with a single value of $p_{V}=0.06$, which is the median value of the three saturnian irregular moons with measured albedos, $0.081\pm 0.02$ for Phoebe (Simonelli et al., 1999), $0.06\pm 0.03$ for Albiorix, and $0.050\pm 0.017$ for Siarnaq (Grav et al., 2015). Substituting in the values for $m_{V,\sun}$ and $p_{V}=0.06$, and turning $r$ into $D$ (diameter), Equation 1 becomes
@@
Table 1: A list of the irregular moons of Saturn with their semi-major axis ($a$), eccentricity ($e$), inclination (ii), orbital period ($P$), V-band absolute magnitude ($H_{V}$), the group they belong to, and the subgroup we have assigned to them (if any). Orbital elements are taken from JPL [^6] https://ssd.jpl.nasa.gov/sats/elem/#\#refs and $H_{V}$ from MPC [^7] https://www.minorplanetcenter.net/iau/NatSats/NaturalSatellites.html. All inclinations are relative to the ecliptic except for Phoebe, which is relative to the Laplace plane.
Name/ $a$ $e$ $i$ $P$ $H_{V}$ Group Subgroup
@@
Albiorix (26) 0.1092 16.33 0.48 36.8 783.5 11.17 Gallic Albiorix
Erriapus (28) 0.1170 17.51 0.476 37.1 871.1 13.71 Gallic Albiorix
@@
Tarvos (21) 0.1218 18.22 0.522 37.8 926.4 13.04 Gallic Albiorix
Bebhionn (37) 0.1138 17.03 0.459 38.5 834.9 14.99 Gallic Albiorix
@@
S/2020 S 3 0.1207 18.06 0.142 46.0 908.2 16.38 Inuit Siarnaq
@@
S/2019 S 6 0.1217 18.21 0.120 46.4 919.7 15.73 Inuit Siarnaq
Siarnaq (29) 0.1195 17.88 0.309 47.8 895.6 10.61 Inuit Siarnaq
Kiviuq (24) 0.0756 11.31 0.275 48.0 449.1 12.67 Inuit Kiviuq
@@
Paaliaq (20) 0.1003 15.00 0.378 48.5 687.0 11.71 Inuit -
Tarqeq (52) 0.1186 17.75 0.143 48.7 885.0 14.82 Inuit Siarnaq
Ijiraq (22) 0.0758 11.34 0.293 49.2 451.4 13.27 Inuit Kiviuq
@@
Bestla (39) 0.1360 20.34 0.486 138.3 1087.2 14.61 Norse -
Narvi (31) 0.1289 19.29 0.441 142.2 1003.9 14.52 Norse -
@@
Hyrrokkin (44) 0.1226 18.34 0.336 149.9 931.9 14.34 Norse -
Skathi (27) 0.1041 15.58 0.281 151.6 728.1 14.41 Norse -
S/2019 S 19 0.1540 23.04 0.458 151.8 1318.0 16.51 Norse Kari
Kari (45) 0.1473 22.03 0.469 153.0 1231.2 14.49 Norse Kari
S/2004 S 21 0.1548 23.16 0.394 153.2 1328.6 16.21 Norse Kari
S/2004 S 36 0.1564 23.39 0.625 153.3 1349.4 16.11 Norse Kari
@@
Geirrod (66) 0.1488 22.26 0.539 154.3 1251.1 15.89 Norse Kari
@@
S/2006 S 1 0.1253 18.75 0.105 156.0 964.2 15.65 Norse Kari
@@
S/2019 S 20 0.1583 23.68 0.354 156.1 1375.5 16.73 Norse Kari
Farbauti (40) 0.1356 20.29 0.249 156.2 1087.3 15.75 Norse Kari
S/2019 S 15 0.1416 21.19 0.257 157.8 1161.6 16.59 Norse Mundilfari
S/2004 S 37 0.1066 15.94 0.448 158.2 755.6 15.92 Norse Mundilfari
@@
Bergelmir (38) 0.1288 19.27 0.145 158.8 1005.5 15.16 Norse Mundilfari
Thiazzi (63) 0.1576 23.58 0.512 158.8 1366.7 15.91 Norse Mundilfari
S/2019 S 5 0.1276 19.09 0.216 158.8 990.4 16.65 Norse Mundilfari
Beli (61) 0.1384 20.70 0.087 158.9 1121.7 16.09 Norse Mundilfari
S/2007 S 9 0.1348 20.17 0.360 158.9 1078.1 16.06 Norse Mundilfari
Skoll (47) 0.1178 17.62 0.463 159.4 878.4 15.41 Norse Mundilfari
@@
Gunnlod (62) 0.1413 21.14 0.251 160.3 1158.0 15.57 Norse Mundilfari
@@
S/2020 S 7 0.1163 17.40 0.500 161.4 861.2 16.79 Norse Mundilfari
S/2020 S 9 0.1699 25.41 0.531 161.4 1532.6 16.02 Norse Mundilfari
@@
Table 2: Continuation of footnote 7.
Name/ $a$ $e$ $i$ $P$ $H_{V}$ Group Subgroup
@@
S/2006 S 13 0.1334 19.95 0.313 162.0 1060.6 16.05 Norse Mundilfari
S/2019 S 16 0.1556 23.27 0.250 162.0 1340.9 16.68 Norse Mundilfari
S/2004 S 53 0.1556 23.28 0.240 162.6 1342.4 16.16 Norse Mundilfari
Jarnsaxa (50) 0.1288 19.27 0.218 163.0 1006.5 15.62 Norse Mundilfari
Gridr (54) 0.1287 19.25 0.187 163.9 1004.7 15.77 Norse Mundilfari
S/2019 S 10 0.1384 20.70 0.248 163.9 1123.0 16.66 Norse Mundilfari
S/2004 S 50 0.1494 22.35 0.45 164.0 1260.4 16.4 Norse Mundilfari
@@
Fenrir (41) 0.1493 22.33 0.137 164.5 1260.2 15.89 Norse Mundilfari
@@
Eggther (59) 0.1326 19.84 0.157 165.0 1052.3 15.39 Norse Mundilfari
S/2004 S 52 0.1768 26.45 0.292 165.3 1633.9 16.5 Norse Mundilfari
Hati (43) 0.1317 19.70 0.372 165.4 1040.1 15.45 Norse Mundilfari
S/2020 S 10 0.1692 25.31 0.296 165.6 1527.2 16.86 Norse Mundilfari
S/2004 S 41 0.1210 18.10 0.301 165.7 914.6 16.31 Norse Mundilfari
S/2004 S 42 0.1219 18.24 0.157 165.7 925.9 16.11 Norse Mundilfari
S/2004 S 39 0.1550 23.19 0.101 165.9 1335.9 16.14 Norse Mundilfari
Aegir (36) 0.1381 20.66 0.255 166.1 1119.3 15.51 Norse Mundilfari
@@
S/2006 S 14 0.1408 21.06 0.060 166.7 1152.7 16.5 Norse Mundilfari
S/2019 S 3 0.1142 17.08 0.249 166.9 837.7 16.22 Norse Mundilfari
S/2020 S 6 0.1420 21.25 0.480 166.9 1167.9 16.55 Norse Mundilfari
Mundilfari (25) 0.1243 18.59 0.212 166.9 952.9 14.57 Norse Mundilfari
S/2019 S 12 0.1396 20.89 0.475 167.1 1138.9 16.33 Norse Mundilfari
S/2004 S 44 0.1305 19.52 0.129 167.7 1026.2 15.82 Norse Mundilfari
S/2004 S 28 0.1462 21.87 0.159 167.9 1220.7 15.77 Norse Mundilfari
S/2004 S 17 0.1317 19.70 0.162 167.9 1040.9 15.95 Norse Mundilfari
@@
Sat LXIV 0.1614 24.14 0.280 168.3 1420.8 16.15 Norse Mundilfari
Surtur (48) 0.1521 22.75 0.448 168.4 1295.6 15.77 Norse Mundilfari
S/2006 S 17 0.1496 22.38 0.425 168.7 1264.5 16.01 Norse Mundilfari
S/2004 S 13 0.1233 18.45 0.265 169.0 942.6 16.25 Norse Mundilfari
S/2004 S 40 0.1075 16.08 0.297 169.2 764.6 16.28 Norse Mundilfari
@@
S/2006 S 18 0.1521 22.76 0.131 169.5 1298.4 16.1 Norse Mundilfari
Fornjot (42) 0.1667 24.94 0.213 170.0 1494.1 15.12 Norse Mundilfari
S/2019 S 4 0.1200 17.95 0.408 170.1 903.9 16.46 Norse Mundilfari
@@
S/2019 S 21 0.1767 26.44 0.155 171.9 1636.4 16.18 Norse Mundilfari
Ymir (19) 0.1534 22.95 0.338 172.3 1315.1 12.41 Norse Phoebe
S/2019 S 8 0.1356 20.29 0.311 172.8 1088.9 16.28 Norse Phoebe
Sat LVIII 0.1745 26.10 0.147 173.0 1604.0 15.7 Norse Phoebe
S/2006 S 9 0.0963 14.41 0.248 173.0 647.9 16.48 Norse Phoebe
@@
S/2007 S 3 0.1311 19.61 0.150 173.8 1034.4 15.74 Norse Phoebe
S/2007 S 2 0.1066 15.94 0.232 174.0 754.9 15.59 Norse Phoebe
S/2006 S 11 0.1318 19.71 0.143 174.1 1042.3 16.47 Norse Phoebe
Greip (51) 0.1229 18.38 0.317 174.1 937.0 15.33 Norse Phoebe
S/2019 S 7 0.1349 20.18 0.232 174.2 1080.6 16.29 Norse Phoebe
Gerd (57) 0.1400 20.95 0.518 174.4 1143.0 15.87 Norse Phoebe
Thrymr (30) 0.1359 20.33 0.467 175.0 1092.2 14.33 Norse Phoebe
@@
S/2019 S 13 0.1401 20.96 0.318 177.3 1144.8 16.68 Norse Phoebe
Alvaldi (65) 0.1470 21.99 0.238 177.4 1232.2 15.62 Norse Phoebe
Angrboda (55) 0.1376 20.59 0.216 177.7 1114.1 16.17 Norse Phoebe